What percent of 150 is 30?

Answer by MRperkins(300)

"is" generally identifies an equals sign.
"of" generally identifies multiplication
so: (what percent x 150 = 30)
or: (n*150=30)
divide each side by 150 and you get n=30/150 or n=1/5 or n=.2
to convert to % you multiply by 100 and you get 20%.
.
private tutoring available
.
contact justin.sheppard.tech@hotmail.com for more information